# The “Little Dark Dot”: Evidence for Self-Interacting Dark Matter in the Strong Lens SDSSJ0946+1006?

### v1

@@ 4.1 Galaxies in the halo
Figure 8: Median stellar-to-halo mass relations from different studies. Colored curves show best-fit relations from P. S. Behroozi et al. (2010), B. P. Moster et al. (2013), P. D. Mitchell et al. (2016), A. Niemiec et al. (2019), G. Girelli et al. (2020), M. Shuntov et al. (2022), and C. Sifón & J. Han (2024), while the TNG-50 simulated result is marked by red dots with $1\sigma$ error bars.
@@
In G. Despali et al. (2024), it is argued that if a dark matter halo with a mass greater than $10^{10}~{}{\rm M_{\odot}}$ exists in J0946+1006, it should already have been detectable in current observations. Although constraining the stellar-to-halo mass relation (SHMR) for SIDM would require full hydrodynamic cosmological simulations, we adopt the reasonable assumption that SIDM follows an SHMR similar to CDM. This allows us to estimate the stellar mass of J0946+1006 if its lensing signal originates from an SIDM halo.
@@
In Fig. 8, we present the SHMR derived from various observational and theoretical studies (P. S. Behroozi et al., 2010; B. P. Moster et al., 2013; P. D. Mitchell et al., 2016; A. Niemiec et al., 2019; G. Girelli et al., 2020; M. Shuntov et al., 2022; C. Sifón & J. Han, 2024). For the TNG-50 simulation (D. Nelson et al., 2019; A. Pillepich et al., 2019), we select halos with masses of $10^{10\pm 0.1}$, $10^{10.5\pm 0.1}$, $10^{11\pm 0.1}$, $10^{11.5\pm 0.1}$, and $10^{12\pm 0.1}~{}{\rm M_{\odot}}$, and calculate the stellar mass distributions of their host galaxies. While the SHMR estimates differ by up to an order of magnitude across different works, they consistently indicate that a dark matter halo with a mass of $10^{11}~{}{\rm M_{\odot}}$ is expected to host a galaxy with a stellar mass in the range of $10^{8}$$10^{9}~{}{\rm M_{\odot}}$. Analysis by G. Despali et al. (2024) based on the Illustris-TNG simulations further supports this result, showing that the mass of the satellite galaxy hosted by the subhalo in J0946+1006 should – given the mass of the subhalo – be larger than $10^{8.5}~{}{\rm M_{\odot}}$.
Observationally, S. Vegetti et al. (2010) estimate the $3\sigma$ detection limit of HST imaging of this lens system to be $5\times 10^{6}~{}{\rm L_{\odot}}$. If a galaxy with a stellar mass of $10^{8}~{}{\rm M_{\odot}}$ remains undetected, it would imply an extraordinarily high mass-to-light ratio of $\sim 20$, which is nearly impossible under normal galaxy formation scenarios. Therefore, while a massive SIDM halo could explain the image perturbations caused by the dark perturber in the lens, the absence of a visible galaxy within it poses a significant new challenge to our understanding of such systems.
4.2 effect of mass-concentration relation
@@
While our fiducial SIDM halos adopt median concentrations based on the $c_{200}(M_{200},z)$ relation from A. A. Dutton & A. V. Macciò (2014), it is necessary to consider scenarios where J0946+1006 significantly exceeds this median value. This is because, if a halo exhibits higher concentration before core-collapse, its core region would accumulate more mass, subsequently leading to a core-collapse central region with greater mass. As a result, lower-mass halos with higher concentrations could produce observational effects similar to those of our fiducial halo. Note that this phenomenon has already been discussed in the background of CDM perturbers (N. C. Amorisco et al., 2022). In E. O. Nadler et al. (2023)’s $N$-body simulation, J0946+1006-like SIDM subhalos also showed enhanced concentrations compared to the field population.
@@
Figure 9: Same format as Fig. 3: dashed curves show the original evolutionary tracks, while solid curves display the evolution of an SIDM halo with pre-collapse $c_{200}$ higher than the median by $3\sigma$ ($=0.45$ dex).
@@
Figure 10: Left & Middle: Same format as Fig. 2, but showing the core-collapsed SIDM halo with color-coded tidal stripping degree. Right: Same format as Fig. 3, but compare the observationally matched SIDM halo across different tidal stripping degree.
The 2D lensing parameters of a high-concentration halo that is undergoing core collapse are compared to those of a fiducial halo in Fig. 9. Observed data are well explained by a subhalo of lower mass ($M_{200}=5\times 10^{10}~{}{\rm M_{\odot}}$), but high concentration ($c_{200}$ above the meidian by $3\sigma$). However, even this lower-mass halo should host a satellite galaxy with mass $5\times 10^{7}~{}{\rm M_{\odot}}$ (according to the SHMR in Fig. 8). For this to be unobservable through emission would still require an unphysical mass-to-light ratio $>$ 10: the mysterious absence of a galaxy in J0946+1006 remains unresolved.
@@ 4.3 Tidal truncation of the dark perturber
So far, we assume the dark halo follows an NFW profile to facilitate comparison with theoretical predictions of the stellar-to-halo mass relation and to estimate the stellar mass of the galaxy it should host. However, if the perturber resides within the host dark matter halo of the lens galaxy, it would inevitably experience tidal stripping. Given that the dark halo required to explain the observational data is highly compact, we estimate that tidal stripping would predominantly remove the halo’s outer regions while leaving its inner profile largely intact.
In left and middle panel of Fig. 10, we remove mass outside different radial ranges of the SIDM halo that best matches the result of Q. Minor et al. (2021) (represented by the red star in Fig. 3). We then compute the 2D projected density profiles and 2D enclosed mass profiles, and compare the lensing properties, as shown in the right panel, with those of the entire SIDM halo. Our theoretical calculations reveal that for a compact core-collapsed SIDM halo, even if more than 95% of the total mass is stripped (corresponding to the mass outside 0.05 $r_{200}$ is removed), $M_{\rm 2D}~{}{\rm(<1~{}kpc)}$ would decrease by merely 4.9%, and $\gamma_{\rm 2D}~{}{\rm(1~{}kpc)}$ would increase by only 9.1%. This indicates that the core properties which produce strong lensing effects remain robust against tidal stripping.
Note that in the above analysis we assume the subhalo has already entered the core-collapse phase before falling into the main halo In reality, if core-collapse occurs after the subhalo falls into the main halo, the environment within the main halo may accelerate the core-collapse process. A detailed analysis of how this process impacts core-collapse would require considering heat conduction, which cannot be handled using the isothermal method. However, in principle, tidal stripping will not affect the mass in the core-collapse region of the halo (H. Nishikawa et al., 2020), and therefore, we do not expect this will impact the main conclusion of this study. E. O. Nadler et al. (2023) also found in their $N$-body simulations that core-collapsed halos are more resilient to tidal disruption by their host halos, further supporting our assumption.
@@ 5 Conclusions
In this work, we conduct a detailed investigation into whether the SIDM scenario can explain the very compact dark perturber previously identified in SDSSJ0946+1006. We find that while a core-collapsed SIDM halo is a viable candidate for the observed perturber, it imposes strict mass requirements. Specifically, only an SIDM halo with $M_{200}\approx 1\times 10^{11}\,{\rm M_{\odot}}$ aligns well with the observational constraints for the perturber in SDSSJ0946+1006. Lower-mass halos, even if they undergo core-collapse, fail to accumulate sufficient mass within the central 1 kpc region to match the observed properties.
@@
By setting the formation time of a $10^{11}\,{\rm M_{\odot}}$ halo equal to the cosmic time, we derive a lower bound on the SIDM cross-section of $\sigma_{m}=138\,{\rm cm^{2}/g}$. This conclusion is largely unaffected by tidal stripping effects on the halo.
@@
Furthermore, we investigate whether fitting an SIDM halo with an NFW-based model could lead to misinterpretations of its mass distribution. We generate a mock lensing image of SDSSJ0946+1006 using the above SIDM profile for the perturber and fit it with a tNFW profile. The resulting best-fit $M_{200}$ is three times lower than the true SIDM halo mass, while the inferred concentration, $c_{200}=62.25$, is significantly overestimated, deviating by $>5\sigma$ from the CDM mass-concentration relation. This outcome is broadly consistent with the findings of Q. Minor et al. (2021).
@@
However, although a core-collapsed halo can explain the observed perturbation, the absence of a visible galaxy within such a massive halo presents a major challenge. For a halo of $1\times 10^{11}~{}{\rm M_{\odot}}$, CDM’s standard stellar-to-halo mass relations predict a central galaxy with $M_{*}\sim 10^{8}-10^{9}~{}{\rm M_{\odot}}$. Even with a $3\sigma$ increase in $c_{200}$ lowering the required SIDM halo mass to $5\times 10^{10}~{}{\rm M_{\odot}}$, a host galaxy with stellar mass $>5\times 10^{7}~{}{\rm M_{\odot}}$ should still have been detectable in existing HST observations. This discrepancy indicates that even a core-collapsed SIDM halo fails to fully account for the observed dark perturber in SDSSJ0946+1006.
@@
Future high-sensitivity follow-up observations with JWST may help determine whether the dark perturber is associated with a galaxy and provide more precise constraints on its mass distribution. Additionally, upcoming space-based surveys, such as those conducted by Euclid, the China Space Station Telescope (CSST), and the Roman Space Telescope, are expected to identify a larger sample of similar perturbers. Such discoveries will be crucial in determining whether compact dark perturbers of this nature are a common phenomenon.
@@
We thank Daneng Yang, Hai-Bo Yu, and Simon White for helpful discussions. SL, RL, KW, XC, and XM acknowledge the support by National Key R&D Program of China No. 2022YFF0503403, the support of the National Nature Science Foundation of China (Nos. 11988101), the support from the Ministry of Science and Technology of China (Nos. 2020SKA0110100), the science research grants from the China Manned Space Project, CAS Project for Young Scientists in Basic Research (No. YSBR-062), and the support from K.C.Wong Education Foundation. ZJ and FJ acknowledge the support of the National Natural Science Foundation of China (Nos. 12473007) and the Beijing Natural Science Foundation (QY23018).
Appendix A Solving the Possion equation
@@
The influence of baryonic matter is not only reflected in the change of the SIDM halo’s density distribution through solving the Poisson equation (Eq. 3), but also the gradual growth of the baryonic potential causes adiabatic contraction in the dark halo, leading to a more centrally concentrated distribution of dark matter (e.g., O. Y. Gnedin et al., 2004). One advantage of the outside-in approach is its flexibility in incorporating baryonic effects on the dark halo’s adiabatic contraction. In the code of F. Jiang et al. (2023), this is achieved by first contracting the CDM halo using the baryonic potential (based on the model of O. Y. Gnedin et al., 2004) to compute $r_{1}$, and then stitching the contracted halo with the inner isothermal core.
@@
To facilitate solving the Poisson equation (Eq. 3) and calculating the adiabatic contraction effect of baryons on the halo, a common approach is to model the baryon density profile using a Hernquist profile (L. Hernquist, 1990):
@@
where $\rho_{\rm b0}$ is determined by the baryon mass $\rho_{\rm b0}=\frac{M_{\rm b}}{2\pi r_{0}^{3}}$ and $r_{0}$ is the scale radius. After substituting the expressions for $\rho_{\rm b}$ and $\rho_{\rm iso}$, Eq. 3 becomes:
@@
For simplicity, we perform the first variable replacement: $x=r/r_{0}$ and $\frac{\mathrm{d}}{\mathrm{d}r}=\frac{1}{r_{0}}\frac{\mathrm{d}}{\mathrm{d}x}$:
@@
Furthermore, we perform the second variable replacement: $y=\frac{x}{1+x}$ and $\frac{\mathrm{d}}{\mathrm{d}x}=(1-y)^{2}\frac{\mathrm{d}}{\mathrm{d}y}$:
@@
Finally by setting $h(y)=-\frac{\Phi(r)}{\nu_{0}^{2}}$ and defining $a\equiv 4\pi Gr_{0}^{2}\rho_{\rm iso0}/\nu_{0}^{2}$ and $b\equiv 4\pi Gr_{0}^{2}\rho_{\rm b0}/\nu_{0}^{2}$, the second-order differential equation can be rewritten in a dimensionless form as:
@@
With the initial conditions $\Phi(0)=0$ and $\Phi^{\prime}(0)=GM_{\rm b}/r_{0}^{2}$ (F. Jiang et al., 2023), this equation can be easily solved numerically.
Appendix B Macro Model Fitting and Results
We utilize the SLaM pipeline to fit the lensing image with the macro model, which ignores the contribution from potential mass perturbations. An overview of the pipeline structure is as follows:
(a) Source Parametric Pipeline: A parametric lens mass and source light model are used to model the lensing image, providing a rapid initial estimation of the lens’s mass. Specifically, the lens mass distribution is described with a Singular Isothermal Ellipsoid (SIE) plus shear model, while the source light is modeled using multiple Gaussian expansion models that include a set of 20 Gaussians sharing a common center and ellipticities.
(b) Source Pixelized Pipeline: Pixelized source models are employed to capture the complex morphology of the background source and refine the lens mass model provided by the Source Parametric Pipeline. Specifically, this pipeline consists of two stages. In the first stage, the lens mass estimation from the Source Parametric Pipeline is inherited as the informative Gaussian prior, in combination with a source pixelization scheme that adapts to the lensing magnification and a constant regularization scheme, to refine the lens mass model and provide an initial estimation of the source’s (potentially irregular) brightness distribution. In the second stage, the lens mass model is fixed to the estimated values given by the first stage, and the pixelized source model is updated to a more advanced one where the pixelization and regularization adapt to the source brightness distribution.
(c) Mass Pipeline: A more general EPL plus shear mass model is used to improve the lens mass estimation given by the Source Pixelized Pipeline. The prior setting for the EPL and shear models are inherited from the best lens mass estimation from the Source Pixelized Pipeline, except for the EPL’s slope, which is assumed to be uniformly distributed between 1.0 and 3.0. For the source model, the pixelized source model from stage 2 of the Source Pixelized Pipeline is used, with all the hyperparameter values fixed.

### v2

@@ 4.1 Galaxies in the halo
Figure 8: Median SHMRs from different studies. Colored curves show best-fit relations from Behroozi2010, Moster2012, Mitchell2015, smhr_Niemiec_2019, Girelli2020, Shuntov_2022, and Sifon2024, while the TNG-50 simulated result is marked by red dots with $1\sigma$ error bars.
@@
In Despali2024, it is argued that if a dark matter halo with a mass greater than $10^{10}~{\rm M_{\odot}}$ exists in J0946+1006, it should already have been detectable in current observations. Although constraining the stellar-to-halo mass relation (SHMR) for SIDM would require full hydrodynamic cosmological simulations, we adopt the reasonable assumption that SIDM follows an SHMR similar to CDM. This allows us to estimate the stellar mass of J0946+1006 if its lensing signal originates from an SIDM halo.
@@
In Fig. 8, we present the SHMR for field halos from various observational and theoretical studies (Behroozi2010; Moster2012; Mitchell2015; smhr_Niemiec_2019; Girelli2020; Shuntov_2022; Sifon2024), a choice motivated by our modeling of the perturber as a complete NFW halo. For the TNG-50 simulation (TNG50-1; TNG50-2), we select halos with masses of $^{10\pm 0.1}, 1010$$^{10.5\pm 0.1}$, $10^{11\pm 0.1}$, $10^{11.5\pm 0.1}$, and $10^{12\pm 0.1}~{\rm M_{\odot}}$ and calculate the stellar mass distributions of their central galaxies. While the SHMR estimates differ by up to an order of magnitude across different works, they consistently indicate that a dark matter halo with a mass of $10^{11}~{\rm M_{\odot}}$ is expected to host a galaxy with a stellar mass in the range of $10^{8}$$10^{9}~{\rm M_{\odot}}$. Analysis by Despali2024 based on the Illustris-TNG simulations further supports this result, showing that the mass of the satellite galaxy hosted by the subhalo in J0946+1006 should – given the mass of the subhalo – be larger than $10^{8.5}~{\rm M_{\odot}}$.
Thus, a massive SIDM halo is expected to host a luminous galaxy, creating a potential tension with the nondetection of a stellar counterpart for the perturber. The initial study by Vegetti_2010 established a stringent $3\sigma$ upper limit of $5\times 10^{6}~{\rm L_{\odot}}$. This limit was derived indirectly by analyzing the residual image – after subtracting a best-fit lens model that assumed a dark subhalo – and measuring the statistical noise within a defined aperture at the perturber’s location. A galaxy with a stellar mass of $10^{8}~{\rm M_{\odot}}$ remaining below this limit would imply an extraordinarily high mass-to-light ratio of $\sim 20$, which is challenging under normal galaxy formation scenarios.
However, the observational constraints on this counterpart are currently uncertain. Recognizing that the residual-based method risks underestimating the brightness, Minor2021 revisited the constraint with a direct modeling approach. They incorporated a Gaussian light profile for the subhalo into their lens model and fitted its luminosity and size simultaneously with all other parameters. This yielded a more conservative upper bound of $1.2\times 10^{8}~{\rm L_{\odot}}$, which the authors cautiously note may be an overestimation due to potential contamination from unsubtracted foreground light.
Moreover, a recent reanalysis of this strong-lensing system by He2025 using PyAutoLens showed that the observed perturbation is better explained by a luminous dwarf satellite with a luminosity of $2.5^{+0.7}_{-0.9}\times 10^{8}\ {\rm L_{\odot}}$ and an NFW halo whose properties are consistent with CDM predictions. However, this statistical preference should be interpreted carefully, as the current observational data are insufficient to unambiguously distinguish this scenario from that of a very compact, purely dark halo.
The significant discrepancy among these observational analyses leaves the perturber’s true luminosity as a pivotal and unresolved issue, thereby preventing a conclusive test of the core-collapse SIDM halo scenario. Ultimately, its viability hinges on future observations sensitive enough to firmly establish the presence or absence of a faint host galaxy.
4.2 Effect of mass–concentration relation
@@
While our fiducial SIDM halos adopt median concentrations based on the $c_{200}(M_{200},z)$ relation from Dutton2014, it is necessary to consider scenarios where J0946+1006 significantly exceeds this median value. This is because, if a halo exhibits higher concentration before core collapse, its core region would accumulate more mass, subsequently leading to a core-collapse central region with greater mass. As a result, lower-mass halos with higher concentrations could produce observational effects similar to those of our fiducial halo. Note that this phenomenon has already been discussed in the background of CDM perturbers (Amorisco2022). In Nadler2023’s $N$-body simulation, J0946+1006-like SIDM subhalos also showed enhanced concentrations compared to the field population.
@@
Figure 9: Same format as Fig. 3: dashed curves show the original evolutionary tracks, while solid curves display the evolution of an SIDM halo with precollapse $c_{200}$ higher than the median by $3\sigma$ (=0.45=0.45 dex).
@@
Figure 10: Left and middle: Same format as Fig. 2, but showing the core-collapse SIDM halo with color-coded tidal stripping degrees. Right: same format as Fig. 3 but comparing the observationally matched SIDM halo across different tidal stripping degrees.
The 2D lensing parameters of a high-concentration halo that is undergoing core collapse are compared to those of a fiducial halo in Fig. 9. Observed data are well explained by a subhalo of lower mass ($M_{200}=5\times 10^{10}~{\rm M_{\odot}}$), but high concentration ($c_{200}$ above the median by $3\sigma$). However, according to the SHMR in Fig. 8, even this lower-mass halo should host a galaxy with mass $5\times 10^{7}~{\rm M_{\odot}}$, and whether this is consistent with observations still remains an unresolved tension due to the current uncertainty in the subhalo’s luminosity.
@@ 4.3 Tidal truncation of the dark perturber
So far, we assume the dark halo follows an NFW profile to facilitate comparison with theoretical predictions of the SHMR and to estimate the stellar mass of the galaxy it should host. However, if the perturber resides within the host dark matter halo of the lens galaxy, it would inevitably experience tidal stripping. Given that the dark halo required to explain the observational data is highly compact, we estimate that tidal stripping would predominantly remove the halo’s outer regions while leaving its inner profile largely intact.
In left and middle panels of Fig. 10, we remove mass outside different radial ranges of the SIDM halo that best matches the result of Minor2021. We then compute the 2D projected density profiles and 2D enclosed mass profiles and compare the lensing properties, as shown in the right panel, with those of the entire SIDM halo. Our theoretical calculations reveal that for a compact core-collapse SIDM halo, even if more than 95% of the total mass is stripped (corresponding to the mass outside 0.05 $r_{200}$ being removed), $M_{\rm 2D}~{\rm(<1~kpc)}$ would decrease by merely 4.9%, and $\gamma_{\rm 2D}~{\rm(1~kpc)}$ would increase by only 9.1%. This indicates that the core properties that produce strong-lensing effects remain robust against tidal stripping.
Note that in the above analysis, we assume the subhalo has already entered the core-collapse phase before falling into the main halo. In reality, if core collapse occurs after the subhalo falls into the main halo, the environment within the main halo may accelerate the core-collapse process. A detailed analysis of how this process impacts core collapse would require considering heat conduction, which cannot be handled using the isothermal method. However, in principle, tidal stripping will not affect the mass in the core-collapse region of the halo (Nishikawa2020); therefore, we do not expect that this will impact the main conclusion of this study. Nadler2023 also found in their $N$-body simulations that core-collapse halos are more resilient to tidal disruption by their host halos, further supporting our assumption.
@@ 5 Conclusions
In this work, we conduct a detailed investigation into whether the SIDM scenario can explain the very compact dark perturber previously identified in SDSS J0946+1006. We find that while a core-collapse SIDM halo is a viable candidate for the observed perturber, it imposes strict mass requirements. Specifically, only an SIDM halo with $M_{200}\approx 1\times 10^{11}\,{\rm M_{\odot}}$ aligns well with the observational constraints for the perturber in SDSS J0946+1006. Lower-mass halos, even if they undergo core collapse, fail to accumulate sufficient mass within the central 1 kpc region to match the observed properties.
@@
By setting the formation time of a $10^{11}\,{\rm M_{\odot}}$ halo equal to the cosmic time, we derive a lower bound on the SIDM cross section of $\sigma_{m}=138\,{\rm cm^{2}/g}$. This conclusion is largely unaffected by tidal stripping effects on the halo.
@@
Furthermore, we investigate whether fitting an SIDM halo with an NFW-based model could lead to misinterpretations of its mass distribution. We generate a mock lensing image of SDSS J0946+1006 using the above SIDM profile for the perturber and fit it with a tNFW profile. The resulting best-fit $M_{200}$ is 3 times lower than the true SIDM halo mass, while the inferred concentration, $c_{200}=62.25$, is significantly overestimated, deviating by $5\sigma$ from the CDM mass–concentration relation. This outcome is broadly consistent with the findings of Minor2021.
@@
However, although a core-collapse SIDM halo can explain the observed perturbation, the absence of a visible galaxy within such a massive halo requires careful consideration of its expected luminosity. For a halo of $1\times 10^{11}~{\rm M_{\odot}}$, CDM’s standard SHMRs predict a central galaxy with $M_{*}\sim 10^{8}-10^{9}~{\rm M_{\odot}}$. Even with a $3\sigma$ increase in $c_{200}$ lowering the required SIDM halo mass to $5\times 10^{10}~{\rm M_{\odot}}$, a host galaxy with a stellar mass $5\times 10^{7}~{\rm M_{\odot}}$ is still expected. This creates a potential tension with the nondetection of a luminous counterpart, an issue that cannot be fully resolved without a definitive measurement of the perturber’s true luminosity.
@@
Future high-sensitivity follow-up observations with the James Webb Space Telescope (JWST) are required to place more precise constraints on the perturber’s mass distribution and definitively determine whether it is associated with a faint host galaxy. Additionally, upcoming space-based surveys, such as those conducted by Euclid, the China Space Station Telescope, and the Roman Space Telescope, will reveal whether such compact dark perturbers are common. This combined approach of detailed characterization and statistical census will be crucial for reshaping our understanding of dark matter.
@@
We thank Daneng Yang, Hai-Bo Yu, Moritz Fischer, and Simon White for helpful discussions. S.L., R.L., K.W., X.C., and X.M. acknowledge the support by National Key R&D Program of China No. 2022YFF0503403, the support of the National Nature Science Foundation of China (No. 11988101), the support from the Ministry of Science and Technology of China (No. 2020SKA0110100), the science research grants from the China Manned Space Project, CAS Project for Young Scientists in Basic Research (No. YSBR-062), and the support from K. C. Wong Education Foundation. Z.J. and F.J. acknowledge the support of the National Natural Science Foundation of China (No. 12473007) and the Beijing Natural Science Foundation (QY23018).
Appendix A Solving the Poisson equation
@@
Not only is the influence of baryonic matter reflected in the change of the SIDM halo’s density distribution through solving the Poisson equation (Eq. 3), but also the gradual growth of the baryonic potential causes adiabatic contraction in the dark halo, leading to a more centrally concentrated distribution of dark matter (e.g., Gnedin2004). One advantage of the outside-in approach is its flexibility in incorporating baryonic effects on the dark halo’s adiabatic contraction. In the code of Jiang2023, this is achieved by first contracting the CDM halo using the baryonic potential (based on the model of Gnedin2004) to compute $r_{1}$ and then stitching the contracted halo with the inner isothermal core.
@@
To facilitate solving the Poisson equation (Eq. 3) and calculating the adiabatic contraction effect of baryons on the halo, a common approach is to model the baryon density profile using a Hernquist profile (Hernquist),
@@
where $\rho_{\rm b0}$ is determined by the baryon mass $\rho_{\rm b0}=\frac{M_{\rm b}}{2\pi r_{0}^{3}}$ and $r_{0}$ is the scale radius. After substituting the expressions for $\rho_{\rm b}$ and $\rho_{\rm iso}$, Eq. 3 becomes
@@
For simplicity, we perform the first variable replacement, $x=r/r_{0}$ and $\frac{\mathrm{d}}{\mathrm{d}r}=\frac{1}{r_{0}}\frac{\mathrm{d}}{\mathrm{d}x}$,
@@
Furthermore, we perform the second variable replacement, $y=\frac{x}{1+x}$ and $\frac{\mathrm{d}}{\mathrm{d}x}=(1-y)^{2}\frac{\mathrm{d}}{\mathrm{d}y}$,
@@
Finally, by setting $h(y)=-\frac{\Phi(r)}{\nu_{0}^{2}}$ and defining $a\equiv 4\pi Gr_{0}^{2}\rho_{\rm iso0}/\nu_{0}^{2}$ and $b\equiv 4\pi Gr_{0}^{2}\rho_{\rm b0}/\nu_{0}^{2}$, the second-order differential equation can be rewritten in a dimensionless form as
@@
With the initial conditions $\Phi(0)=0$ and $\Phi^{\prime}(0)=GM_{\rm b}/r_{0}^{2}$ (Jiang2023), this equation can be easily solved numerically.
Appendix B Macro Model Fitting
We utilize the SLaM pipeline to fit the lensing image with the macro model, which ignores the contribution from potential mass perturbations. An overview of the pipeline structure is as follows.
(a) Source parametric pipeline. A parametric lens mass and source light model are used to model the lensing image, providing a rapid initial estimation of the lens’s mass. Specifically, the lens mass distribution is described with a singular isothermal ellipsoid plus shear model, while the source light is modeled using multiple Gaussian expansion models that include a set of 20 Gaussians sharing a common center and ellipticities.
(b) Source pixelized pipeline. Pixelized source models are employed to capture the complex morphology of the background source and refine the lens mass model provided by the source parametric pipeline. Specifically, this pipeline consists of two stages. In the first stage, the lens mass estimation from the source parametric pipeline is inherited as the informative Gaussian prior, in combination with a source pixelization scheme that adapts to the lensing magnification and a constant regularization scheme, to refine the lens mass model and provide an initial estimation of the source’s (potentially irregular) brightness distribution. In the second stage, the lens mass model is fixed to the estimated values given by the first stage, and the pixelized source model is updated to a more advanced one, where the pixelization and regularization adapt to the source brightness distribution.
(c) Mass pipeline. A more general EPL plus shear mass model is used to improve the lens mass estimation given by the source pixelized pipeline. The prior settings for the EPL and shear models are inherited from the best lens mass estimation from the source pixelized pipeline, except for the EPL’s slope, which is assumed to be uniformly distributed between 1.0 and 3.0. For the source model, the pixelized source model from stage 2 of the source pixelized pipeline is used, with all the hyperparameter values fixed.